Exercise 15.38 Enhanced with Feedback You may want to reference (D ages 693-702) section 15.8 while completing this problem Consider the following reaction: H2 (g) I2 (g) 2 HI (g) Complete the following table. Assume that all concentrations are equilibrium concentrations in M. CHIl Kel 0.922 0.0355 0.0388 4.60x10-2 M 0.385 M 90.6 340 445 4.90x102 M 4.64x10-2 M 50.2 Part A Find KC at 25 o C. Express your answer using three significant figures. Submit My Answers Give Up Part B Find [H2] at 340 GC. Express your answer using two significant figures Submit My Answers Give Up Part C Find DHIJ at 445 °C. Express your answer using three significant figures. CHI] Submit My Answers Give Up M. M.
Answer #1

Ans. Balanced reaction:       H2(g) + I2(g) ------------> 2 HI(g)

            Equilibrium constant, Kc = [HI]2 / ([H2] [I2])                  - equation 1

Part A: Putting the values in equation 1-

Kc = (0.922)2/ (0.0355 x0.0388) = 0.850084 / 0.0013774 = 617.16

Hence, Kc = 617

Part B. Putting the values in equation 1-

90.6 = (0.385)2/ ([H2] x 4.6 x10-2)

Or, 90.6 x 0.046 [H2] = 0.148225

Or, [H2] = 0.148225 / 4.1676 = 0.04

Thus, [H2] = 0.04 M

Part C: Putting the values in equation 1-

50.2 = ([HI])2/ (4.9 x 10-2 x 4.64 x10-2)

Or, ([HI])2 = 50.2 x 0.0022736 = 0.11413472

Or, [HI] = (0.11413472)1/2 or under root (0.11413472)

Hence, [HI] = 0.34

Thus, [HI] = 0.34